OpenStudy (baldymcgee6):

Would somebody mind checking my work? The equation x^2y+2xy^3=8 defines y as a function of x, y=f(x), near x=2, y =1. Find the slope of the curve x^2y+2xy^3=8 when x=2, y=1

OpenStudy (baldymcgee6):

Here is my work

OpenStudy (anonymous):

Double check the denominator of the final line.

OpenStudy (baldymcgee6):

oops, thank you for noticing that, forgot to square the first 2 in the denominator. is everything else right?

OpenStudy (anonymous):

I think so. I'm going through it one more time to be sure.

OpenStudy (anonymous):

-3/8.

OpenStudy (baldymcgee6):

Sorry, meant to type -6/16 or -3/8

OpenStudy (anonymous):

Everything else is fine. Final line: \[\large \frac{-2(2)(1)-2(1)^3}{2^2+6(2)(1)^2}=\frac{-3}{8}\]

OpenStudy (baldymcgee6):

Thanks so much @CliffSedge

OpenStudy (anonymous):

Hey, my pleasure. That was all very clear and easy to check!